Consider the following income statement for the Heir Jordan Corporation: HEIR JORDAN CORPORATION Income Statement Sales Costs $43,800 34,800 Taxable income Taxes (35%) $ 9,000 3,150 Net income 5,850 Dividends Addition to retained earnings S 3,300 2,550 The projected sales growth rate is 12 percent Prepare a pro forma income statement assuming costs vary with sales and the dividend payout ratio is constant. (Input all amounts as positive values. Do not round intermediate calculations.) HEIR JORDAN CORPORATION Pro Forma Income Statement Sales Costs Taxable income Taxes Net income What is the projected addition to retained earnings? (Do not round intermediate calculations.) Addition to retained earnings
Answer #1

From given data,

Projected Sales Growth Rate = 12%

Sales = 1.12* 43800

Sales = $49,056

Costs = (34800/43800)*49056

Costs = $38,976

Taxable Income = 49056 - 38976

Taxable Income = $10,080

Taxes = 0.35 * 10080

Taxes = $3,528

Net Income = 10080 - 3528

Net Income = $6,552

Addition to Retained Earnings = (2550/5850) * 6552

Addition to Retained Earnings = $2,856